OpenStudy (anonymous):

Let F = (−2x + sin(yz)) i  + (4y + sin(xz)) j  + (3z + sin(xy)) k  Let S be the cube of side length 2 centered at (4, 0, 0) and with faces parallel to the coordinate planes. Let S1 be the face of S which is perpendicular to the x-axis and whose x values are greater than 4. (a) Find the flux of F through S1, oriented in the positive x-direction. (b) Find the flux of F outward through the remaining parts of the cube (all sides except S1).

OpenStudy (anonymous):

is there a special case or do i have to use the long general formula

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

you need to work part a by setting up a flux integral you can use divergence thm for part b

OpenStudy (anonymous):

\[\int\limits \int\limits (-2x+\sin(yz))i +(4y+\sin(xz))j + (-(1/3)\sin(xy))k * (+2i-4j+k)dxdy\]

OpenStudy (anonymous):

that would be the integral for part a? @ganeshie8

OpenStudy (anonymous):

what would my limits of integration be? 0 to 4 for dx and 0to2 for dy?

ganeshie8 (ganeshie8):

can you try -32 for part a once ?

OpenStudy (anonymous):

no I don't have tries, it was due Tuesday but I want to understand this material

OpenStudy (anonymous):

how did you get -32? did i set up the integral right? @ganeshie8

ganeshie8 (ganeshie8):

whats the normal vector of S1 ?

OpenStudy (anonymous):

I don't know?

OpenStudy (anonymous):

the x value is greater than 4 but I don't know what the normal is @ganeshie8

OpenStudy (anonymous):

so do you think my integral is wrong for part a?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

how did u get 2i-4j+k ?

ganeshie8 (ganeshie8):

you need to find the normal vector of surface to find the flux through the surface

OpenStudy (anonymous):

i use the flux integral general formula

OpenStudy (anonymous):

2i+4j+k is (-f_xi - f_yj +k)

OpenStudy (anonymous):

@ganeshie8 is the normal just <4,0,0>

ganeshie8 (ganeshie8):

that gives you normal vector of surface but why are you messing with your Force field ?

ganeshie8 (ganeshie8):

Right! the normal is just <4, 0, 0>

ganeshie8 (ganeshie8):

the unit nromal is just <1, 0, 0>

OpenStudy (anonymous):

i was trying to use the general formula of flux integral

ganeshie8 (ganeshie8):

okay parameterize the surface

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\int\limits_{?}^{?} F (x,y, f(x,y)) * (-f_x-f_y+k)dxdy\]

OpenStudy (anonymous):

this is the formula i was trying to use, where f(x,y) is z= -(1/3) sin(x,y)

OpenStudy (anonymous):

@ganeshie8 this formula does not work here?

OpenStudy (anonymous):

\[x=(-2+sinyz))+4t ....... y=(4y+\sin(xz))+0t.......... z+(3z+\sin(xy))+0t\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

that works always but it would be ridankulous to use it for S1

ganeshie8 (ganeshie8):

your surface is very simply, you're told that it is perpendicular to x axis so the normal vector is <1,0,0>

ganeshie8 (ganeshie8):

you can use below parameterization : x = 4 y = y z = z

OpenStudy (anonymous):

so that's just doing way to much work?

ganeshie8 (ganeshie8):

\[\int\limits_{?}^{?}\int\limits_{?}^{?} F (4,y, z) * (1,0,0)dydz\]

ganeshie8 (ganeshie8):

setup the bounds and evaluate ^

OpenStudy (anonymous):

so if the cube is lenght 2 and center at (4,0,0) dy is 0 to 2 and dz is 0 to 2 ? @ganeshie8

OpenStudy (anonymous):

why is it " * (1,0,0)" ?

OpenStudy (anonymous):

@ganeshie8 for part b is it the divF * (volume of the cube) - part a?

ganeshie8 (ganeshie8):

\[\hat{n}dS = \langle 1,0,0\rangle dydz \]

ganeshie8 (ganeshie8):

you're almost correct about bounds

OpenStudy (anonymous):

so we don't need the normal <4,0,0>?

OpenStudy (anonymous):

so is it -1 to 1?

ganeshie8 (ganeshie8):

http://gyazo.com/22c04d8d064622a3b5ec4843246bb11a

ganeshie8 (ganeshie8):

Yes! both y and z bounds would be from -1 to +1

ganeshie8 (ganeshie8):

\[\int\limits_{-1}^{1}\int\limits_{-1}^{1} F (4,y, z) * (1,0,0)dydz \]

ganeshie8 (ganeshie8):

\[\int\limits_{-1}^{1}\int\limits_{-1}^{1} (-2(4) + \sin(yz))dydz \]

ganeshie8 (ganeshie8):

evaluate ^

ganeshie8 (ganeshie8):

thats for part a

ganeshie8 (ganeshie8):

you're correct about partb just find the entire flux out of cube using divergence thm and subtract the part a

OpenStudy (anonymous):

what would dA be?

OpenStudy (anonymous):

curlF = _2i +4j+3k volume of cube= 8

OpenStudy (anonymous):

@ganeshie8 so 5* 8 +32

OpenStudy (anonymous):

72?

ganeshie8 (ganeshie8):

Looks good !

OpenStudy (anonymous):

thank you so much @ganeshie8